# Multigraph Message Passing with Bi-Directional Multi-Edge Aggregations

### v1

@@ Abstract
Graph Neural Networks (GNNs) have seen significant advances in recent years, yet their application to multigraphs, where parallel edges exist between the same pair of nodes, remains under-explored. Standard GNNs, designed for simple graphs, compute node representations by combining all connected edges at once, without distinguishing between edges from different neighbors. There are some GNN architectures proposed specifically for multigraph tasks, yet these architectures perform only node-level aggregation in their message-passing layers, which limits their expressive power. Furthermore, these approaches either lack permutation equivariance when a strict total edge ordering is absent, or fail to preserve the topological structure of the multigraph. To address all these shortcomings, we propose MEGA-GNN, a unified framework for message passing on multigraphs that can effectively perform diverse graph learning tasks. Our approach introduces a two-stage aggregation process in the message passing layers: first, parallel edges are aggregated, followed by a node-level aggregation that operates on aggregated messages from distinct neighbors. We show that MEGA-GNN supports permutation equivariance and invariance properties. We also show that MEGA-GNN is universal given a strict total order on the edges. Experiments on synthetic and real-world financial transaction datasets demonstrate that MEGA-GNN either significantly outperforms or is on par with the accuracy of state-of-the-art solutions.
@@ 1 Introduction
Our theoretical analyses show that MEGA-GNN is provably powerful, meaning it can detect any directed subgraph pattern within multi-graphs if a strict total ordering of the edges is possible.
Our main contributions are three fold: (1) We introduce a generic bi-directional message passing framework for multigraphs that incorporates artificial nodes between pairs of nodes to aggregate parallel edges before message passing, which makes it effective across diverse multigraph learning tasks. (2) We prove that our framework supports essential properties such as permutation equivariance and universality. (3) We validate our framework on financial transaction datasets, covering detection of illicit transactions and phishing accounts. We outperform the state-of-the-art for illicit transaction detection and match the state-of-the-art results for phishing account detection.
@@ 2 Limitations of the existing solutions
Multi-GNN introduced a provably powerful GNN architecture with simple adaptations for directed multigraphs. A notable contribution of Multi-GNN was the multigraph port ID assignment on edges, which allows the model to differentiate between edges from the same neighbor and edges from different neighbors, an essential feature for handling multigraph structures (see Figure 1). In addition to multigraph port numbering, Multi-GNN incorporates ego-IDs (You et al. (2021)) and bi-directional message passing (Jaume et al. (2019)). The authors also proved that with these three adaptations, it is possible to assign unique node IDs in connected directed multigraphs, making their solution universal.
The main drawback of Multi-GNN is that augmenting edge features with multigraph port numbers leads to the loss of permutation invariance and equivariance properties (see Proposition 2.1 and the proof provided in Appendix A.1). These properties are critical for many graph learning tasks to ensure that the model’s predictions remain consistent under arbitrary node or edge permutations.
@@
The multigraph port numbering scheme of Egressy et al. (2024), is not permutation-equivariant in the absence of a contextually-driven strict total ordering of edges.
@@
ADAMM proposed aggregating all parallel edges between two nodes to a single undirected super-edge, thereby transforming the multigraph into a simple graph before message passing. An initial features for this super-edge is computed using DeepSet and the subsequent message passing layers operate on this features. However, this approach results in a loss of critical structural information inherent in the multigraph. Specifically, by collapsing multiple edges into a single one, the formulation can not generate features for individual edges. Consequently, it cannot perform graph learning tasks that require edge features.
Notably, because this approach does not compute features for the original edges, it cannot perform multi-edge aggregations repeatedly in several message passing layers (see Figure 2).
@@
In summary, a general framework for message passing on multigraphs that supports permutation invariance and a general set of graph learning tasks has yet to be proposed. This paper addresses this gap by introducing a novel message passing architecture that is explicitly designed for multigraphs.
@@ 3 Multigraph Message Passing
This section introduces a novel message-passing scheme for multigraphs which combines node-level aggregation with multi-edge aggregation and supports a general set of graph learning tasks such as node, edge, and graph classification. Additionally, the resulting models are permutation invariant or equivariant depending on the chosen multi-edge aggregation function and the downstream task.
@@ 3.1 Motivation for Multi-Edge Aggregations in Multigraph GNNs
We argue that a two-stage aggregation approach is more powerful than a single-stage aggregation in a multigraph setting, wherein aggregation of the parallel edges between the same source and destination nodes is performed before a node-level aggregation. We first give a motivating example, which shows that two-stage aggregation functions are more expressive than their single-stage counterparts.
Figure 3: Comparison of two-stage and single-stage aggregation strategies for multigraphs. The left panel illustrates a sample multi-graph. The center panel depicts the two-stage aggregation process, wherein the first stage aggregates parallel edges, followed by a second stage node-level aggregation. The right panel shows the single-stage aggregation scheme commonly employed by standard GNNs, which would aggregate all edges, including the parallel edges, at once.
Consider the example graph depicted in Figure 3, with four edges $e_{1},e_{2},e_{3}$ and $e_{4}$. Suppose that we have an aggregation function that operates on multisets and can simultaneously compute both the sum and the maximum of the elements included in the set. To demonstrate that two-stage aggregation approach is more powerful than the single-stage one, we will show that two-stage approach can perform any computation that the single-stage approach can, but not vice versa.
A single-stage scheme, can compute only $(\text{SUM}\{\!\!\{e_{1},e_{2},e_{3},e_{4}\}\!\!\}\ \text{and}\ \text{MAX}\{\!\!\{e_{1},e_{2},e_{3},e_{4}\}\!\!\})$. The two-stage approach can also compute these results by aggregating the edges in pairs first: $(\text{SUM}\{\!\!\{\text{SUM}\{\!\!\{e_{1},e_{2}\}\!\!\},\text{SUM}\{\!\!\{e_{3},e_{4}\}\!\!\}\}\!\!\},\text{MAX}\{\!\!\{\text{MAX}\{\!\!\{e_{1},e_{2}\}\!\!\},\text{MAX}\{\!\!\{e_{3},e_{4}\}\!\!\}\}\!\!\})$. However, the two-stage aggregation allows for more nuanced computations that the single-stage approach cannot perform. For example, the two-stage approach can compute $\text{SUM}\{\!\!\{\text{MAX}\{\!\!\{e_{1},e_{2}\}\!\!\},\text{MAX}\{\!\!\{e_{3},e_{4}\}\!\!\}\}\!\!\}$ and $\text{MAX}\{\!\!\{\text{SUM}\{\!\!\{e_{1},e_{2}\}\!\!\},\text{SUM}\{\!\!\{e_{3},e_{4}\}\!\!\}\}\!\!\}$, which the single-stage approach cannot compute.
This distinction is significant; the two-stage aggregation enables us to capture the statistics of edges originating from different neighbors. Consider a financial application, in which a receiver may receive the highest total payment from one sender (Sender 1), while the highest single payment may originate from another sender (Sender 2). Using the two-stage scheme, we can detect the maximum of the sums of payments from each sender, allowing us to distinguish between their contributions. In contrast, the single-stage scheme can only compute the maximum among all payments without differentiating between identical and distinct senders, which can obscure critical insights.
@@ 3.2 Preliminaries
In this work, we consider attributed graphs. The nodes and the edges of the graph ${\mathcal{G}}$ are associated with initial feature vectors $\mathbf{x}_{i}^{(0)}\in\mathbb{R}^{D_{n}}$ and $\mathbf{e}_{ij}^{(0)}\in\mathbb{R}^{D_{e}}$, and latent feature vectors $\mathbf{x}_{i}^{(l)}\in\mathbb{R}^{D^{h}_{n}}$ and $\mathbf{e}_{ij}^{(l)}\in\mathbb{R}^{D^{h}_{e}}$, where $D_{n},D_{e},D^{h}_{n}$ and $D^{h}_{e}\in\mathbb{R}$ represent the dimensions of the given vectors, respectively, and $l\in[1,L]$, $L$ is the number of layers of the model.
In a multigraph, there can be parallel edges between the same two nodes, each with its own feature vector, which requires additional considerations. Let ${\mathcal{E}}^{supp}\subseteq{\mathcal{E}}$ denote the support set of ${\mathcal{E}}$, which consists of unique $(i,j)$ pairs that have a multiplicity of at least one in the multi-set ${\mathcal{E}}$. Furthermore, let $ME_{ij}=\{\!\{(i,j)\in{\mathcal{E}}\}\!\}$ denote the multiset of parallel edges between a source node $i$ and a destination node $j$. Assume also that the cardinality of $ME_{ij}$ is $P_{ij}$. The initial and the latent feature vectors of these parallel edges are denoted with $\mathbf{e}_{ijp}^{(0)}\in\mathbb{R}^{D_{e}}$ and $\mathbf{e}_{ijp}^{(l)}\in\mathbb{R}^{D^{h}_{e}}$, where $p\in\{1..P_{ij}\}$.
@@
$\begin{split}\mathbf{x}_{j}^{(l)}=g_{v}^{(l-1)}\bigg{(}\mathbf{x}_{j}^{(l-1)},\text{AGG}\{\!\{f^{(l-1)}(\mathbf{x}_{i}^{(l-1)},\mathbf{e}_{ijp}^{(l-1)})\ |\ i\in N_{in}(j),p\in\{1..P_{ij}\}\ \}\!\}\bigg{)},\end{split}$ (1)
where $f^{(l)}$ is a function that constructs the messages from the incoming neighbors of node $i$. These messages are then aggregated using the AGG function, which is typically a permutation-invariant operation such as sum, mean, or max. After the message aggregation step, the latent features of node $i$ is updated using the update function $g_{v}^{(l)}$. Additionally, latent edge features could also be updated using the following formula, where $g_{e}^{(l)}$ is the edge update function:
@@
(Strict Total Order) A multiset of edges ${\mathcal{E}}$ in a graph is said to have a strict total order under a binary relation $<$ if the following conditions hold for every $a$, $b$ and $c$in ${\mathcal{E}}$: (1) $a\neq b$, either $a<b$ or $b<a$, (2) not $a<a$, (3) if $a<b$ and $b<c$, then $a<c$ (see Munkres (2000), p. 22).
@@ 3.3 Multigraph Message Passing with Multi-Edge Aggregations
Commonly used GNN architectures Kipf & Welling (2017); Veličković et al. (2018); Hamilton et al. (2017) are not designed for multigraphs. Performing message aggregation solely at the node level as shown in Equation 1, can limit the expressive power of these models on multigraph learning tasks. As a remedy, we propose a two-stage aggregation process, wherein the first stage, termed multi-edge aggregation, aggregates parallel edges, followed by a second stage node-level aggregation. To enable effective multi-edge aggregation, we introduce novel artificial nodes into the multigraph.
We define the set of artificial nodes, positioned between each unique pair of nodes in the multigraph as shown in Figure 4 (b) as follows:
@@
The incoming neighbors of the artificial node $v_{ij}^{art}$ is given by the multiset $ME_{ij}$, the cardinality of which is $P_{ij}$. Let $\mathbf{h}_{ij}\in\mathbb{R}^{d}$ be the latent features of the artificial node $v_{ij}^{art}$.
The message-passing algorithm for multigraphs is adapted to perform information aggregation in two stages. First, parallel edges are aggregated at the artificial nodes, thereby reducing the multigraph to an equivalent simple graph:
$\begin{split}\mathbf{h}_{ij}^{(l-1)}=\text{EdgeAgg}\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in\{1..P_{ij}\}\}.\end{split}$ (4)
In the second stage, aggregation occurs at the original graph nodes:
$\begin{split}\mathbf{a}_{j}^{(l-1)}&=\text{AGG}\{f^{(l-1)}(\mathbf{x}_{i}^{(l-1)},\mathbf{h}_{ij}^{(l-1)})\ |\ i\in N_{in}(j)\}\\ \mathbf{x}_{j}^{(l)}&=g_{v}^{(l-1)}\bigg{(}\mathbf{x}_{j}^{(l-1)},\mathbf{a}_{j}^{(l-1)}\bigg{)},\end{split}$ (5)
where $N_{in}(j)=\{i\in{\mathcal{V}}|(i,j)\in{\mathcal{E}}^{supp}\}$, which is not a multiset, it only includes distinct incoming neighbors of node $j$. As in 2, the latent features of each individual edge can be updated without loss of generality:
@@
This two-stage aggregation process, enabled by our innovative use of artificial nodes, preserves the topology of the original multigraph. Equations 4, 5, and 6 can be applied iteratively, allowing for effective propagation and updates of both node and edge latent features in each message-passing layer. Unlike previous methods such as ADAMM Sotiropoulos et al. (2023), which collapses parallel edges into a single edge prior to message passing—thereby preventing updates on the original multigraph edges—our approach maintains distinct edges through individual updates, enabling richer feature propagation in multigraph learning tasks.

### v2

@@ Abstract
Graph Neural Networks (GNNs) have seen significant advances in recent years, yet their application to multigraphs, where parallel edges exist between the same pair of nodes, remains under-explored. Standard GNNs, designed for simple graphs, compute node representations by combining all connected edges at once, without distinguishing between edges from different neighbors. There are some GNN architectures proposed specifically for multigraphs, yet these architectures perform only node-level aggregation in their message passing layers, which limits their expressive power. Furthermore, these approaches either lack permutation equivariance when a strict total edge ordering is absent, or fail to preserve the topological structure of the multigraph. To address all these shortcomings, we propose MEGA-GNN, a unified framework for message passing on multigraphs that can effectively perform diverse graph learning tasks. Our approach introduces a two-stage aggregation process in the message passing layers: first, parallel edges are aggregated, followed by a node-level aggregation of messages from distinct neighbors. We show that MEGA-GNN is not only permutation equivariant but also universal given a strict total ordering on the edges. Experiments show that MEGA-GNN significantly outperforms state-of-the-art solutions by up to 13% on Anti-Money Laundering datasets and is on par with their accuracy on real-world phishing classification datasets in terms of minority class F1 score.
@@ 1 Introduction
Our theoretical analyses show that MEGA-GNN is provably powerful, meaning it can detect any directed subgraph pattern within multigraphs if a strict total ordering of the edges is possible.
Our main contributions are three fold: (1) We introduce a generic bi-directional message passing framework for multigraphs using two-stage aggregation where parallel edges are first aggregated, followed by a node-level aggregation of messages from distinct neighbors in each message passing layer. (2) We prove that our framework supports essential properties such as permutation equivariance and universality. (3) We validate our framework on financial transaction datasets, where we significantly outperform the state-of-the-art for illicit transaction detection and match the state-of-the-art for phishing account detection.
@@ 2 Limitations of the existing solutions
Multi-GNN introduced a provably powerful GNN architecture with simple adaptations for directed multigraphs. A notable contribution of Multi-GNN is the multigraph port numbering on edges, which allows the model to differentiate between edges from the same neighbor and edges from different neighbors (see Figure 1). In addition to multigraph port numbering, Multi-GNN incorporates Ego-IDs (You et al. (2021)) and bi-directional message passing (Jaume et al. (2019)). The authors also proved that with these three adaptations, it is possible to assign unique node IDs in connected directed multigraphs, making their solution universal.
The main drawback of Multi-GNN is that augmenting edge features with multigraph port numbers leads to the loss of permutation equivariance (see Proposition 2.1 and the proof provided in Appendix A.1). Permutation equivariance is critical for many graph learning tasks to ensure that the model’s predictions remain consistent under arbitrary node or edge permutations.
@@
The multigraph port numbering scheme of Egressy et al. (2024), is not permutation equivariant in the absence of a contextually-driven strict total ordering of edges.
@@
ADAMM proposed aggregating all parallel edges between two nodes to a single undirected super-edge, thereby transforming the multigraph into a simple graph before message passing. An initial features for this super-edge are computed using DeepSet (Zaheer et al. (2017)) and the subsequent message passing layers operate on these features. However, this approach results in a loss of critical structural information inherent in the multigraph. By collapsing multiple edges into single one, it cannot generate features for individual edges, limiting its ability to perform graph learning tasks that require latent edge features.
Notably, because this approach does not compute latent features for the original edges, it cannot perform multi-edge aggregations repeatedly in several message passing layers.
@@
In summary, a general framework for message passing on multigraphs that supports permutation equivariance and diverse graph learning tasks has yet to be proposed. This paper addresses this gap by introducing a novel message passing architecture that is explicitly designed for multigraphs.
@@ 3 Multigraph Message Passing
This section introduces a novel message passing scheme for multigraphs which combines node-level aggregation with multi-edge aggregation and supports a general set of graph learning tasks such as node, edge, and graph classification. Additionally, the resulting models are permutation invariant or equivariant depending on the chosen multi-edge aggregation function and the downstream task.
@@ 3.1 Motivation for Multi-Edge Aggregations in Multigraph GNNs
Figure 3: Comparison of two-stage and single-stage aggregation strategies for multigraphs. The left panel illustrates a sample multigraph. The center panel depicts the two-stage aggregation process, wherein the first stage aggregates parallel edges, followed by a second stage node-level aggregation. The right panel shows the single-stage node-level aggregation scheme commonly employed by standard GNNs, which aggregates all edges, including the parallel edges, at once.
We argue that a two-stage aggregation is more powerful than a single-stage aggregation in a multigraph setting. We first give a motivating example, which shows that two-stage aggregation functions are more expressive than their single-stage counterparts.
Consider the example graph depicted in Figure 3, with four edges $e_{ij1},e_{ij2},e_{kj1}$ and $e_{kj2}$. Suppose that we have an aggregation function that operates on multisets and can simultaneously compute the sum and the maximum of the elements included in the multiset. To demonstrate that two-stage aggregation is more powerful than the single-stage one, we will show that two-stage aggregation can perform any computation that the single-stage aggregation can, but not vice versa. A single-stage aggregation can compute the following:
$\text{SUM}\{e_{ij1},e_{ij2},e_{kj1},e_{kj2}\},\quad\text{MAX}\{e_{ij1},e_{ij2},e_{kj1},e_{kj2}\}.$
The two-stage aggregation can also compute these by aggregating the edges in pairs first:
$\text{SUM}\{\text{SUM}\{e_{ij1},e_{ij2}\},\text{SUM}\{e_{kj1},e_{kj2}\}\},\quad\text{MAX}\{\text{MAX}\{e_{ij1},e_{ij2}\},\text{MAX}\{e_{kj1},e_{kj2}\}\}.$
On the other hand, the two-stage aggregation allows for more nuanced computations, such as:
$\text{SUM}\{\text{MAX}\{e_{ij1},e_{ij2}\},\text{MAX}\{e_{kj1},e_{kj2}\}\},\quad\text{MAX}\{\text{SUM}\{e_{ij1},e_{ij2}\},\text{SUM}\{e_{kj1},e_{kj2}\}\},$
which cannot be achieved through single-stage aggregation. The distinction is significant; the two-stage aggregation enables us to capture the statistics of edges originating from different neighbors. Consider a financial application, in which a receiver may receive the highest total payment from one sender (Sender 1), while the highest single payment may originate from another sender (Sender 2). Using the two-stage aggregation, we can detect the maximum of the sums of payments from each sender, allowing us to distinguish between their total contributions. In contrast, the single-stage aggregation can only compute the maximum among all payments without differentiating between identical and distinct senders, which can obscure critical insights.
@@ 3.2 Preliminaries
Let ${\mathcal{E}}^{supp}\subseteq{\mathcal{E}}$ denote the support set of ${\mathcal{E}}$, which consists of unique $(i,j)$ pairs that have a multiplicity of at least one in the multiset ${\mathcal{E}}$. Furthermore, let $ME_{ij}=\{\!\{(i,j)\in{\mathcal{E}}\}\!\}$ denote the multiset of parallel edges between a source node $i$ and a destination node $j$. We denote the cardinality of $ME_{ij}$ as $P_{ij}$. For a node $j\in{\mathcal{V}}$, we define the incoming and outgoing neighbors as $N_{in}(j)=\{i\in{\mathcal{V}}|(i,j)\in{\mathcal{E}}^{supp}\}$ and $N_{out}(j)=\{i\in{\mathcal{V}}|(j,i)\in{\mathcal{E}}^{supp}\}$, respectively.
The nodes and edges of the graph ${\mathcal{G}}$ are associated with feature vectors. Each node and edge has an initial feature vector $\mathbf{x}_{i}^{(0)}\in\mathbb{R}^{D_{n}}$ and $\mathbf{e}_{ijp}^{(0)}\in\mathbb{R}^{D_{e}}$, with $p\in[1..P_{ij}]$, respectively. Similarly, at layer $l\in[1..L]$ in a model with $L$ layers, nodes and edges are represented by latent feature vectors $\mathbf{x}_{i}^{(l)}\in\mathbb{R}^{D^{h}_{n}}$ and $\mathbf{e}_{ijp}^{(l)}\in\mathbb{R}^{D^{h}_{e}}$, respectively. Here $D_{n},D_{e},D^{h}_{n},D^{h}_{e}\in\mathbb{N}$ denote the dimensions of the corresponding vectors.
@@
$\begin{split}\mathbf{x}_{j}^{(l)}=g_{v}^{(l-1)}\bigg{(}\mathbf{x}_{j}^{(l-1)},\text{AGG}\{\!\{f^{(l-1)}(\mathbf{x}_{i}^{(l-1)},\mathbf{e}_{ijp}^{(l-1)})\ |\ i\in N_{in}(j),p\in[1..P_{ij}]\ \}\!\}\bigg{)},\end{split}$ (1)
where $f^{(l-1)}$ is a function that constructs the messages from the incoming neighbors of node $j$. These messages are then aggregated using the AGG function, which is typically a permutation-invariant operation such as sum, mean, or max. After the message aggregation step, the latent features of node $j$ is updated using the update function $g_{v}^{(l-1)}$. Additionally, latent edge features could also be updated using the following formula, where $g_{e}^{(l-1)}$ is the edge update function:
@@
(Strict Total Order) Strict total order is a binary relation $<$ on a set ${\mathcal{A}}$, which satisfies the following conditions for all $a$,$b$,$c$in ${\mathcal{A}}$: (1) $a\neq b$, either $a<b$ or $b<a$, (2) not $a<a$, (3) if $a<b$ and $b<c$, then $a<c$ (see Munkres (2000), p. 22).
We define a strict total order on the edges of the attributed graph ${\mathcal{G}}$ using their initial feature vectors. Specifically, for any two edges, the ordering is established by: $\mathbf{e}_{ijp}^{(0)}<\mathbf{e}_{i^{\prime}j^{\prime}q^{\prime}}^{(0)}$.
@@ 3.3 Multigraph Message Passing with Multi-Edge Aggregations
Commonly used GNN architectures Kipf & Welling (2017); Veličković et al. (2018); Hamilton et al. (2017) are not designed for multigraphs. Performing message aggregation solely at the node level as shown in Equation 1, limits the expressive power of these models on multigraph learning tasks. As a remedy, we propose a message passing framework (MEGA-GNN) for multigraphs with two-stage aggregation, wherein the first stage, termed multi-edge aggregation, aggregates parallel edges, followed by a second stage node-level aggregation. To enable effective multi-edge aggregation, we introduce novel artificial nodes into the multigraph.
We define the set of artificial nodes, positioned between each unique pair of nodes in the multigraph (as shown in Figure 4(b)) as follows:
@@
The incoming neighbors of the artificial node $v_{ij}^{art}$ are given by the multiset $ME_{ij}$, the cardinality of which is $P_{ij}$. Let $\mathbf{h}_{ij}\in\mathbb{R}^{d}$ be the latent features of the artificial node $v_{ij}^{art}$.
In the first aggregation stage, parallel edges are aggregated at the artificial nodes, thereby reducing the multigraph to an equivalent simple graph in the message passing layer $(l\!-\!1)$:
$\begin{split}\mathbf{h}_{ij}^{(l-1)}=\text{EdgeAgg}\{\!\{\mathbf{e}_{ijp}^{(l-1)}\ |\ p\in[1..P_{ij}]\}\!\}.\end{split}$ (4)
In the second aggregation stage, node-level aggregation is performed.
$\begin{split}\mathbf{a}_{j}^{(l-1)}&=\text{AGG}\{\!\{f^{(l-1)}(\mathbf{x}_{i}^{(l-1)},\mathbf{h}_{ij}^{(l-1)})\ |\ i\in N_{in}(j)\}\!\}\\ \mathbf{x}_{j}^{(l)}&=g_{v}^{(l-1)}\bigg{(}\mathbf{x}_{j}^{(l-1)},\mathbf{a}_{j}^{(l-1)}\bigg{)}.\end{split}$ (5)
Here, the destination node $j$ receives a single (aggregated) message from each of its distinct incoming neighbors. As in Equation 2, the latent features of each individual edge can be updated:
@@
This two-stage aggregation process, enabled by our innovative use of artificial nodes, preserves the topology of the original multigraph. Equations 4, 5, and 6 can be applied iteratively, allowing for effective propagation and updates of both node and edge latent features in each message passing layer. Unlike previous methods such as ADAMM Sotiropoulos et al. (2023), which collapses parallel edges into a single edge prior to message passing—thereby preventing updates on the original multigraph edges—our approach maintains distinct edges through individual updates, enabling richer feature propagation in multigraph learning tasks.